OpenStudy (anonymous):

What are the solutions -3-y^2=24?

OpenStudy (anonymous):

\[Only 4 answers. A: is \pm3\sqrt{3} B: \pm3i \sqrt{3} C: \pm9\sqrt{3} D: \pm9i \sqrt{3}\]

OpenStudy (anonymous):

only one of those four answers.